OpenStudy (anonymous):

The yearly attendance at a local restaurant is 54,200 and grows continuously at a rate of 7.2% each year. What is the approximate attendance at the restaurant in 21 years? 136,150 58,239 245,840 233,394

jimthompson5910 (jim_thompson5910):

Use the formula A = Pe^(r*t) where in this case P = 54200 r = 0.072 t = 21

OpenStudy (anonymous):

but the answer in not in my answer choices

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

81950.4

jimthompson5910 (jim_thompson5910):

show me what you typed in

OpenStudy (anonymous):

54200(0.072*21)

jimthompson5910 (jim_thompson5910):

you forgot about the e

jimthompson5910 (jim_thompson5910):

the formula is A = Pe^(r*t)

OpenStudy (anonymous):

whats e ?

jimthompson5910 (jim_thompson5910):

e is like pi, but in this case e = 2.71828182846....

jimthompson5910 (jim_thompson5910):

your calculator should know what to do with 'e' when you type it in

jimthompson5910 (jim_thompson5910):

if not, then just use that number for 'e'

jimthompson5910 (jim_thompson5910):

so you should have 54200*e^(0.072*21) or 54200*2.71828182846^(0.072*21)